OpenStudy (anonymous):

There are 45 people working in a particular building. A survey on dietary habits is conducted by giving every third person in the building a questionnaire to fill out. From the questionnaire the researcher obtains a score on that person’s dietary habits. What is the number of members in the population?

OpenStudy (studygurl14):

every third, meaning 1 out of 3 people given the questionnaire. (aka \(\large\frac{1}{3}\)) so...\(\large45\times\frac{1}{3}\rightarrow\frac{45}{3}\rightarrow15\)

OpenStudy (studygurl14):

to give me a medal, click best answer